Independent solution

How to solve this Conditional Probability question

Setup

Setup

Relate the supplied intersection to the conditional accident rate within the male group.

Pr⁡(M∩A)=Pr⁡(A∣M)Pr⁡(M)\Pr(M\cap A)=\Pr(A\mid M)\Pr(M)

Model

Model

Solve the product identity for the male marginal probability.

Pr⁡(M)=Pr⁡(M∩A)Pr⁡(A∣M)=0.300.50=0.60\Pr(M)=\frac{\Pr(M\cap A)}{\Pr(A\mid M)}=\frac{0.30}{0.50}=0.60

Compute

Compute

The two driver categories are complementary.

Pr⁡(F)=1−Pr⁡(M)=1−0.60=0.40\Pr(F)=1-\Pr(M)=1-0.60=0.40

Answer

Answer

The female proportion is 0.40.

0.40(B)\boxed{0.40\quad\text{(B)}}